Consider an economy that is represented by: C= 8,000+0.75Yd I= 4,000-400i G= 6,000 T=0.2Y L=0.2Y-100i M=60000 P=20, Compute for the values of equilibrium levels of output and interest rate
1
Expert's answer
2019-10-15T06:27:15-0400

Y=C+I+G-T, so Y=8,000+0.75Y+4,000-400i+6,000-0.2Y, after simplifying 0.45Y=18,000-400i. From the other side M/P=L, so 60000/20=0.2Y-100i, after simplifying 3000=0.2Y-100i. After solving the system of equations Y=24,000, i=1.8